# ShortGPT: Layers in Large Language Models are More Redundant Than You Expect

### 06 Mar 2024 v1

@@ 6 Related works
Model pruning: model pruning LeCun et al. (1989); Han et al. (2015) is a classic and effective method of reducing model redundancy modules to compress models. The model pruning methods mainly include unstructured pruning and structured pruning. The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT Frantar and Alistarh (2023) and LoRAPrune Zhang et al. (2023). However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition. Another more practical approach is structured pruning, GUMSyed et al. (2023) makes an analysis of several structured pruning methods for decoder-only LLMs. LLM-Pruner Ma et al. (2024) selectively removes non-critical structures according to gradient information. ShearedLLaMA Xia et al. (2023) employs targeted structured pruning and dynamic batch loading. LaCo Yang et al. (2024) used layer merging to compress the model. Compared to the previous method, our method is a simple and efficient structured pruning method.
@@
Knowledge distillation: Knowledge distillation (KD) Hinton et al. (2015); Gou et al. (2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network. Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results Li et al. (2022). In addition, when we can access transparent teacher models, the result can be improved more. MiniLLM Gu et al. (2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution. DistlLLM Ko et al. (2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs. However, compared to model pruning, this method often requires higher computational resources.
@@
Quantization: quantization Liu et al. (2021); Gholami et al. (2022); Dettmers et al. (2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models. Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms. LUT-GEMM Park et al. (2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format. SPQR Dettmers et al. (2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits. Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
@@
Low-rank factorization: low-rank decomposition Cheng et al. (2017); Povey et al. (2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions. TensorGPT Xu et al. (2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices. Recently, SliceGPT Ashkboos et al. (2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
@@
Model redundancy: researchers have long noticed the significant redundancy in nonlinear models Catchpole and Morgan (1997). In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy. In Bian et al. (2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads. In Dalvi et al. (2020), researchers dissect two pre-trained models, BERT Devlin et al. (2018) and XLNet Yang et al. (2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level. However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
@@ Appendix C Layer Removal on Baichuan2-series Model
We also found that layer removal has a significant impact on the performance of the model in generative tasks. Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval Chen et al. (2021), which is a dataset used to evaluate the performance of code generation models. The results are shown in Figure 10. After removing 10 layers, the performance on HumanEval almost completely degraded. We speculate that it may be due to significant cumulative errors in generative tasks.

### 07 Mar 2024 v2

@@ 6 Related works
Model pruning: model pruning (LeCun et al., 1989; Han et al., 2015) is a classic and effective method of reducing model redundancy modules to compress models. The model pruning methods mainly include unstructured pruning and structured pruning. The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar and Alistarh, 2023) and LoRAPrune (Zhang et al., 2023). However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition. Another more practical approach is structured pruning, GUM(Syed et al., 2023) makes an analysis of several structured pruning methods for decoder-only LLMs. LLM-Pruner (Ma et al., 2024) selectively removes non-critical structures according to gradient information. ShearedLLaMA (Xia et al., 2023) employs targeted structured pruning and dynamic batch loading. LaCo (Yang et al., 2024) used layer merging to compress the model. Compared to the previous method, our method is a simple and efficient structured pruning method.
@@
Knowledge distillation: Knowledge distillation (KD) (Hinton et al., 2015; Gou et al., 2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network. Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results (Li et al., 2022). In addition, when we can access transparent teacher models, the result can be improved more. MiniLLM (Gu et al., 2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution. DistlLLM (Ko et al., 2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs. However, compared to model pruning, this method often requires higher computational resources.
@@
Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models. Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms. LUT-GEMM (Park et al., 2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format. SPQR (Dettmers et al., 2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits. Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
@@
Low-rank factorization: low-rank decomposition (Cheng et al., 2017; Povey et al., 2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions. TensorGPT (Xu et al., 2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices. Recently, SliceGPT (Ashkboos et al., 2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
@@
Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole and Morgan, 1997). In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy. In (Bian et al., 2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads. In (Dalvi et al., 2020), researchers dissect two pre-trained models, BERT (Devlin et al., 2018) and XLNet (Yang et al., 2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level. However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
@@ Appendix C Layer Removal on Baichuan2-series Model
We also found that layer removal has a significant impact on the performance of the model in generative tasks. Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval (Chen et al., 2021), which is a dataset used to evaluate the performance of code generation models. The results are shown in Figure 10. After removing 10 layers, the performance on HumanEval almost completely degraded. We speculate that it may be due to significant cumulative errors in generative tasks.